OpenStudy (anonymous):

Simplify; cot^2 x - csc^2 x

OpenStudy (anonymous):

This is actually a demoninator, but I already simplified the numerator which was originaly "cos^2 x + sin^2 x" into 1.

OpenStudy (anonymous):

it is pythagorean, much like the first identity you described... it would be -1

OpenStudy (anonymous):

Oh, I see. I don't really need an explanation but I'm interested into how you got that.

OpenStudy (jdoe0001):

see the last formulat on that picture? solve by \(\large cot^2(\theta)\)

OpenStudy (anonymous):

I remember my pythagorean's by taking the original cos^2(x)+sin^2(x)=1 and dividing each term by either of the terms... such as, divide each term by cos^2(x)... leads to 1+tan^2(x)=sec^2(x).... the other one you get by dividing each term by sin^2(x).

OpenStudy (anonymous):

@jdoe I don't know what to do with the cot ._.

OpenStudy (anonymous):

That's right @jdoe0001, just like that.

OpenStudy (anonymous):

I guess this just goes to show I've got plenty of tutoring sessions to attend. Thanks you two!

OpenStudy (anonymous):

You're welcome.

OpenStudy (jdoe0001):

$$ 1+cot^2(\theta)=\color{blue}{csc^2(\theta)} \implies cot^2(\theta)=\color{blue}{csc^2(\theta)}-1 $$